A bond with a face value of $ 1,000.00, an 18% coupon that pays interest annually, matures in 10 years. If your required rate of return is 12%

How much would you be willing to pay for the bonus today?
Answer #1

Price of the bond = PV of cash flows from it

Period Cash flows PVF/PVAF @6% Disc cash flows
1-20 $      90.00 11.4699 $        1,032.29
20 $1,000.00 0.3118 $           311.80
Price of the Bond $        1,344.10

PVF = 1/ (1+r)^n

r = Interest rate

n = Time gap

PVAF = SUm [ PVF(r%, n) ]
PVF(r%, n) = 1 / ( 1 + r)^n
r = Int rate per period
n = No. of periods

Semi annualy interest paid , so Interest = 1000*18%*6/12

= 90

PVF/PVAF shoulls be 6% as it is semi annual

= 12% * 6/12

= 6%

Pls do rate, if the answer is correct and comment, if any further assistance is required.